# Quantum Long Short-Term Memory (QLSTM) vs Classical LSTM in Time Series Forecasting: A Comparative Study in Solar Power Forecasting

### v2

@@ 6.2 Performance Analysis
The analysis of generalization performance, illustrated in Table 7, demonstrates the QLSTM model’s superiority in terms of generalization, substantiated by lower mean and median test loss values over all epochs compared to the classical LSTM model. This performance, coupled with more accurate and reliable predictions, holds the potential to revolutionize solar power production forecasting. Figure 6 provides a comparative analysis of the train and test loss for both models.
@@ 9 Conclusion
Moreover, our controlled experiments unveiled a consistent pattern – QLSTMs outperformed classical LSTMs in minimizing test loss across multiple metrics, including MAE, MSE, and RMSE. This elevated forecasting precision, coupled with heightened generalization capabilities, positions QLSTMs as a disruptive force in the renewable energy sector. By empowering grid operators with unparalleled foresight into solar supply fluctuations, QLSTMs could catalyze more judicious energy management, storage optimization, and seamless incorporation of photovoltaic sources into existing infrastructure.
@@
Conflict of Interest
The authors declare no conflict of interest or competing interests
Data Availability
Data available upon request.
Authors’ Contribution
All authors contributed equally to this work.
@@ A.1 LSTM
Figure 7: LSTM Circuit 9747369
@@ A.2 QLSTM
The QLSTM replaces LSTM cells with 6 variational quantum circuits (VQCs) to form a quantum LSTM cell. VQCs leverage a small number of qubits and gates to represent complex functions. This quantum layer showed quicker convergence and more stable loss than the classical LSTM 9747369 . The information flow in a quantum LSTM cell is described by the equations:
@@
Figure 8: Generic VQC architecture for QLSTM. It consists of three layers: the data encoding layer (with the H, Ry, and Rz gates), the variational layer (dashed box), and the quantum measurement layer. 9747369
Figure 9: QLSTM Circuit 9747369

### v3

@@ 6.2 Performance Analysis
The analysis of generalization performance, illustrated in Table 7, demonstrates the QLSTM model’s superiority in terms of generalization, substantiated by lower mean and median test loss values over all epochs compared to the classical LSTM model. This performance, coupled with more accurate and reliable predictions, holds the potential to enhance solar power production forecasting. Figure 6 provides a comparative analysis of the train and test loss for both models.
@@
The right subfigure in Figure 3 shows the test loss over epochs for both QLSTM and classical LSTM models. The classical LSTM model (red curve) demonstrates a generally good fit, with test loss decreasing initially and then stabilizing with minor fluctuations. This indicates that the classical LSTM is effectively learning the data and maintaining reasonable generalization.
However, the QLSTM model (blue curve) exhibits a more stable and lower test loss across epochs, suggesting that it generalizes marginally better than the classical LSTM. The consistent performance of QLSTM, with less fluctuation in test loss, highlights its ability to capture complex patterns while maintaining robustness in prediction accuracy. This slight edge in generalization makes QLSTM a promising alternative for solar power forecasting.
@@
6.3 Evaluation Time
The evaluation time per epoch is a critical factor in assessing the practicality of the Quantum Long Short-Term Memory (QLSTM) and classical Long Short-Term Memory (LSTM) models, particularly for applications requiring rapid predictions. The QLSTM model, which leverages quantum computational principles, exhibits significantly longer evaluation times due to the complex nature of quantum simulations. On average, each epoch of the QLSTM model required approximately 5172.22 seconds (approximately 1 hour and 26 minutes). This extended duration is attributed to the intensive quantum computations and the simulation environment provided by Pennylane.
In contrast, the classical LSTM model, which operates within a conventional computational framework, demonstrated a much shorter average evaluation time of approximately 0.41 seconds per epoch. This stark difference in evaluation times underscores the trade-off between the advanced capabilities of quantum models and the computational efficiency of classical models. While QLSTM achieves superior accuracy and faster convergence, its longer evaluation time may limit its applicability in scenarios where rapid predictions are essential.
Model Time (seconds)
QLSTM 5172.22
LSTM 0.41
Table 8: Comparison of per epoch average evaluation times
@@ 9 Conclusions
Moreover, our controlled experiments unveiled a consistent pattern, QLSTMs outperformed classical LSTMs in minimizing test loss across multiple metrics, including MAE, MSE, and RMSE. This elevated forecasting precision, coupled with heightened generalization capabilities, positions QLSTMs as a disruptive force in the renewable energy sector. By empowering grid operators with unparalleled foresight into solar supply fluctuations, QLSTMs could catalyze more judicious energy management, storage optimization, and seamless incorporation of photovoltaic sources into existing infrastructure.
@@
Conflict of interest/Competing interests
The authors declare no conflict of interest or competing interests.
Availability of data and materials
Data is available upon request.
@@ A.1 LSTM
Figure 7: LSTM Circuit [8]
@@ A.2 QLSTM
The QLSTM replaces LSTM cells with 6 variational quantum circuits (VQCs) to form a quantum LSTM cell. VQCs leverage a small number of qubits and gates to represent complex functions. This quantum layer showed quicker convergence and more stable loss than the classical LSTM [8]. The information flow in a quantum LSTM cell is described by the equations:
@@
Figure 8: Generic VQC architecture for QLSTM. It consists of three layers: the data encoding layer (with the H, Ry, and Rz gates), the variational layer (dashed box), and the quantum measurement layer. [8]
Figure 9: QLSTM Circuit [8]